OpenStudy (legacy0110):

Find the range of the function: F(x)=(x-2)^2+2

OpenStudy (marcelie):

range lowest y values to highest y values

OpenStudy (marcelie):

aah this is tricky hmm im thinking that it might be (2 , infinity )

OpenStudy (marcelie):

https://www.desmos.com/calculator

OpenStudy (sshayer):

\[\left( x-2 \right)^2\ge0\] \[\left( x-2 \right)^2+2 \geq0+2\] \[\left( x+2 \right)^2+2 \geq 2\]

OpenStudy (sshayer):

correction it is (x-2)^2

OpenStudy (marcelie):